OpenStudy (anonymous):

Find the slope of the line that contains the following points. A(5, 6), B(10, 8) 5/2 2/5 14/15

OpenStudy (cmadaviss):

it should be be 2/5. You can use desmos calculator to put the two points in.

OpenStudy (09395548052cs):

Using the formula to solve the slope with a given points, m= y sub 2- y sub 1/ x sub 2 - x sub 1 m=8-6/10-5 =2/5